# M2: An Analytic System with Specialized Storage Engines for Multi-Model Workloads

### v1

@@ 1. Introduction
Unlike traditional workloads based on a single data model, modern data analytic workloads increasingly require the ability to handle multiple data models simultaneously. In the fields of machine learning and artificial intelligence, in particular, workloads typically involve processing array data for matrix computations or linear algebra algorithms together with structured and semi-structured data. Consider, for example, an e-commerce recommendation system that builds a non-negative matrix factorization model (Lee and Seung, 2000). In this scenario, order and review data, both stored in a document-oriented model, are combined to construct a recommendation matrix that represents rating values for products by customers. This matrix is then fed into the factorization algorithm to generate recommendations. To extract insights such as identifying products that a customer would like, the reconstructed matrix is converted into a relational format for further analysis using relation operations. This data analysis pipeline requires handling three distinct data models, namely, relational, document-oriented, and array models. There are many use cases reported in the literature for combining an array data model with structured and semi-structured data models (Koren et al., 2009; Ferdousi et al., 2017; Xian et al., 2020).
@@ 2.2. Interface and Operation
Listing 1 demonstrates how M2 can be used to predict the top 10 favorable products for a given customer during targeted marketing campaigns. The query begins by accessing order and review collections which are stored in a document-oriented model (Line 1). A document in the order collection includes a customer identifier, while a review document contains identifiers for an order and a product, and a rating value. To obtain rating values for each customer and product, these collections are merged using the join function, followed by the project function to filter only relevant dimensions and rating values (Line 2). Before being fed into the non-negative matrix factorization algorithm, the set of rating values is converted into a matrix by the toArray function with names for dimensions and a value (Line 3). By using the number of tuples in customer and product tables and the rand function, the factor matrices W and H are initialized (Lines 5-7), and the algorithm updates these matrices (Lines 8-10). To focus the analysis on specific customer-product pairs of interest (for targeted marketing campaigns), the interest table containing customer and product identifiers is opened (Line 12). Once the recommendation matrix is reconstructed (Line 13), the join function takes interest which is in a relational model and filled in an array model, producing equi-joined relational data as an output (this inter-model join operation will be discussed later). Among the joined rating values, the ones with the customer id = 3 are selected, and the top ten items are extracted (Line 14). Finally, the execute function is called to get results from the described query above (Line 16).
@@ 3.1. Query Execution Flow
When a user submits a multi-model query through M2’s querying interface, the system processes it through several stages. First, the query planner interprets the query, performing binding operations to create a logical plan represented as a directed acyclic graph (DAG). This logical plan consists of nodes and edges representing operations and producer-consumer relationships between nodes, respectively, capturing all operations across different data models and their relationships. The plan is then partitioned based on the data models involved. Each partition contains operations specific to a particular model (relational, document, or array) that will be executed by the corresponding storage engine. DuckDB handles both relational and document operations 111 We enhanced DuckDB with VelocyPack (ArangoDB, Inc., 2025b), a JSON serialization format, and added specialized JSON processing operations to optimize document processing in DuckDB. , while PreVision processes array computations. A bridge module sits between the execution engines, handling inter-model conversions and join operations.
@@ 4. Multi-Stage Hash Join
Algorithm 1 describes the MSHJ procedure between array and relational data models, taking data from both models as input. Let denote the array size, the tile size, and the number of dimensions. For clarity, we use a relation as the non-array input model, but a join between a collection and an array can be performed in the same way, as documents in collections can also contain dimension attributes.
@@ 5. Evaluation
We conducted an evaluation to assess the effectiveness of M2. All our evaluations were conducted on a desktop machine with an Intel i7-9700K CPU, 32GB RAM, and a 1TB SSD running Linux. All systems used in evaluations were configured to utilize 24GB of memory for their buffer spaces. Before each experiment, we reset both the system-specific buffer spaces and the operating system’s page cache to eliminate any performance advantages from cached data. We configured the systems to utilize a single thread to see clearer performance characteristics of each storage engine. For systems using multiple storage engines, we configured them to avoid interleaving query executions and to use the same amount of buffer space.

### v2

@@ 1. Introduction
Unlike traditional workloads based on a single data model, modern data analytic workloads increasingly require the ability to handle multiple data models simultaneously. In the fields of machine learning and artificial intelligence, in particular, workloads typically involve processing array data for matrix computations or linear algebra algorithms together with structured and semi-structured data. Consider, for example, an e-commerce recommendation system that builds a non-negative matrix factorization model. In this scenario, order and review data, both stored in a document-oriented model, are combined to construct a recommendation matrix that represents rating values for products by customers. This matrix is then fed into the factorization algorithm to generate recommendations. To extract insights such as identifying products that a customer would like, the reconstructed matrix is converted into a relational format for further analysis using relation operations. This data analysis pipeline requires handling three distinct data models, namely, relational, document-oriented, and array models. There are many use cases reported in the literature for combining an array data model with structured and semi-structured data models (Koren et al., 2009; Ferdousi et al., 2017; Xian et al., 2020).
@@ 2.2. Interface and Operation
Listing 1 demonstrates how M2 can be used to predict the top 10 favorable products for a given customer during targeted marketing campaigns. The query begins by accessing order and review collections which are stored in a document-oriented model (Line 1). A document in the order collection includes a customer identifier (cid), while a review document contains a product identifier (pid) and a rating value. To obtain rating values for each customer and product, these collections are joined on order identifiers (oid) using the join function, followed by the project function to filter only relevant dimensions and rating values (Line 2). Before being fed into the non-negative matrix factorization algorithm, the set of rating values is converted into a matrix by the toArray function with names for dimensions and a value (Line 3). By using the number of tuples in customer and product tables and the rand function, the factor matrices W and H are initialized (Lines 5-7), and the algorithm updates these matrices (Lines 8-10). To focus the analysis on specific customer-product pairs of interest (for targeted marketing campaigns), the interest table containing customer and product identifiers is opened (Line 12). Once the recommendation matrix is reconstructed (Line 13), the join function takes interest which is in a relational model and filled in an array model, producing equi-joined relational data as an output (this inter-model join operation will be discussed later). Among the joined rating values, the ones with the cid = 3 are selected, and the top ten items are extracted (Line 14). Finally, the execute function is called to get results from the described query above (Line 16).
@@ 3.1. Query Execution Flow
When a user submits a multi-model query through M2’s querying interface, the system processes it through several stages. First, the query planner interprets the query, performing binding operations to create a logical plan represented as a directed acyclic graph (DAG). This logical plan consists of nodes representing operations and edges indicating data flow relationships between producer and consumer operations, capturing all operations across different data models and their relationships. The plan is then partitioned based on the data models involved. Each partition contains operations specific to a particular model (relational, document, or array) that will be executed by the corresponding storage engine. DuckDB handles both relational and document operations 111 We enhanced DuckDB with VelocyPack (ArangoDB, Inc., 2025b), a JSON serialization format, and added specialized JSON processing operations to optimize document processing in DuckDB. , while PreVision processes array computations. A bridge module sits between the execution engines, handling inter-model conversions and join operations.
@@ 4. Multi-Stage Hash Join
Algorithm 1 describes the MSHJ procedure between array and relational data models, taking data from both models as input. Let denote the array size, the tile size, and the number of dimensions. For clarity, we use a relation as the non-array input model, but a join between a collection and an array can be performed in the same way, as documents in collections can also contain dimension attributes. We use 0-based indexing throughout this section.
@@ 5. Evaluation
We conducted an evaluation to assess the effectiveness of M2. All our evaluations were conducted on a desktop machine with an Intel i7-9700K CPU, 32GB RAM, and a 1TB SSD running Linux. All systems used in evaluations were configured to utilize 24GB of memory for their buffer spaces. Before each experiment, we reset both the system-specific buffer spaces and the operating system’s page cache to eliminate any performance advantages from cached data. We configured the systems to utilize a single thread to see clearer performance characteristics of each storage engine. For systems using multiple storage engines, we configured them to avoid interleaving query executions and to use the same amount of buffer space for each data model.